# A quantitative analysis of Galilei’s observations of Jupiter satellites from the Sidereus Nuncius

### v1

@@ Abstract
We analyse the observations of the satellites of Jupiter from the Sidereus Nuncius (January 7 to March 1, 1610) and compare them to the predictions obtained using a modern sky simulator, verifying them one by one. A sinusoidal fit of the data obtained from the 64 available sketches, allows measuring the relative major semi-axes of the satellites’ orbits and their periods with a statistical precision of 2-4% and 0.1-0.3% respectively. The periods are basically unbiased while the orbits tend to be underestimated for Callisto by about 12%. The posterior fit error indicates that the positions of the satellites are determined with a resolution of 0.4-0.6 Jupiter diameters in the notation of Galilei corresponding to about 40-70 arc sec i.e. similar to the true angular diameter of Jupiter, in those days. We show that with this data one can infer in a convincing way the third law of Kepler for the Jupiter system. The 1:2 and 1:4 orbital resonance between the periods of Io and Europa/Ganymede can be determined with % precision. In order to obtain these results it is important to separate the four datasets. This operation, which is nowadays simple using a sky simulator, and is fully reported in this work, was an extremely difficult task for Galilei as the analysis will evidence. Nevertheless we show how the four periods might have been extracted using the modern Lomb-Scargle technique without having to separate the four data-sets already just using these early observations. We also perform a critical evaluation of the accuracy of the observation of the Pleiades and other clusters and the Moon. To conclude we use these data to infer the possible resolution and aberration of the “cannocchiale” of Galilei and use these parameters to imagine what could have been the view offered by Galilei to the Doge when he was promoting his instrument from the Campanile of San Marco in Venice. This analysis allows quantifying and appreciating the level of accuracy that Galilei managed to attain with the instrumentation he had at his disposal.
@@ 1 Introduction
The night of January 7 1610 was a special one for Galilei and, in hindsight, for the history of science. With an improved version [^1] Interestingly, Galilei writes that he had not seen the new stars with the previous version of the telescope. of his telescope, that he had already advertised about six months earlier to the headquarters of the Venetian republic, Galilei points to Jupiter.
He sees three peculiar brilliant “stars” very close to the planet and strangely well aligned to the Ecliptic plane (not a random one!). See Fig. 22. Probably by false modesty, he writes he had not paid too much attention to the relative distances as he expected them to be fixed stars and, guided by I know not what fate, (nescio quo fato ducto) he came back to observe it the night after. It is hard to believe he did not already had an intuition of the exceptional discovery that was behind the corner. Which is the probability of having by change three aligned stars so closely packed to Jupiter? And in that very meaningful direction? Quite small indeed. But something even more amazing was happening the day after.
@@
On January 8 Jupiter appears to the East of the three stars (Fig. 23). In reality the easternmost “star” (Callisto) was still there to the East of Jupiter, but Galilei does not see it, unlikely to the earlier night. It is interesting to notice that Callisto will always be observed in the following two months, unless when it was too close to Jupiter. The three stars were then in reality not the same three satellites: Io had passed from the East to the West side and Europa had become visible to the West after being too close to Jupiter the day before.
If the stars were fixed the explanation could be that Jupiter was moving in the “wrong” direction (eastward) in the sky. Galilei thinks that Jupiter, as well as Mars, might be travelling in retrograde motion. Clouds on the 9th create some frustration but a new observation becomes possible on the 10th: Jupiter is now West of two stars now!
@@
A thrilling adventure that will accompany him uninterruptedly for each clear night for the following two months was just about to start [^2] It is remarkable how many clear days there were in Padua during those winter months: just ten days were cloudy out of about sixty. On average nowadays the fraction of cloudy days is 40-50% according to [1].. This detailed account of the dance of the satellites of Jupiter is an history of scientific rigour and excitement: it lead him to observe until late at night in several occasions despite the many duties that he must have had in Padua.
Earlier in the book Galilei writes that astronomers should engage in determining the periods of the satellites as this was not possible to him due to the restrictions in time. In view of this consideration it is amazing how accurately Galilei decided to document this early phenomenon. Despite his statement these 64 observations were indeed extremely rich and go far beyond what one would deem sufficient to state a new “discovery”.
@@
It is possible that these observations seemed so exceptional and groundbreaking to him to deserve such a solid and complete basis of evidence despite this might seem redundant to our eyes.
@@
It should also be not forgotten that Galilei might have felt the pressure of releasing the book as soon as possible due to the risk of being anticipated by other scientists, which was not completely an unlikely scenario. Despite this he deemed necessary to report 64 observations. This choice tells a lot about of the mindset of Galilei and this emphasis on data cannot but induce a certain degree of admiration in modern experimentalists.
In this work we analyse the observations of the satellites of Jupiter reported by Galileo Galilei in the Sidereus Nuncius [2] (“The stellar messenger”) during the period from January 7 to March 1, 1610. The 64 sketches that Galilei presented in his revolutionary work are presented in Sec. 2 and the result of the image digitization in Sec. 3. In Sec. 3.1 we convert the time of the observations from “Italian hours”, calculated starting from sunset, into modern hours (CET) to match it with the predictions of an online sky simulator [3] “Stellarium” (Sec. 3.4). Before performing the association of the measurements to the individual four satellites, we discuss the possibility of determining the orbital parameters from a global analysis of all the points taken together (Sec. 3.3): we consider the statistical distribution of the satellites position and an analysis based on the Lomb-Scargle periodogram in Sec. 3.3.
While performing the association of the satellites (Sec. 3.4)) we highlight in detail the possible ambiguities due to the fact that sometimes a pair of close satellites were reported as that they could be missed due to their proximity to the planet’s disk. We also show how, generally, the observation have a clear mapping to the expectation from the simulation.
A sinusoidal fit of the data obtained from the data-sets divided by satellite, is presented in Sec. 3.6. We discuss the achievable accuracy on the determination of the major semi-axes of the satellites’ orbits and their periods and the residuals between the observations and the fit. We also consider the origin of the observed biases with respect to modern determinations. We show that with this data one can infer in a convincing way the equivalent of the third law of Kepler for the Jupiter system. We show that the 1:2 and 1:4 orbital resonances between the periods of Io : Europa and Io : Ganymede can be determined with % precision from these data.
In Sec. 4 we expand the study by performing a critical evaluation of the accuracy of the observation of the Pleiades cluster, the Beehive cluster, the region of Orion belts, what Galilei called the ”head of Orion Nebula” and finally the Moon.
To conclude in Sec. 3.7 we use the accuracy of Jupiter data to infer the possible resolution and aberration of the “cannocchiale” of Galilei. We use these parameters to imagine what could have been the view offered by Galilei to the head of the republic of Venice (the “Doge” of the “Serenissima”) when he was promoting his instrument from the Campanile of San Marco in Venice.
2 Available dataset from the “Stellar messenger”
@@
In this work we decided to only focus on the observations presented in the Sidereus Nuncius without considering for simplicity later observations performed in later years by Galileo himself. The images were taken from [2]. We show in Figs. 1-2 a collection of the sketches of Sidereus Nuncius in chronological order with a progressive numeric labeling.
@@
In total Galileo took 64 sketches in 1610 from January 7th up to March 1st basically inspecting Jupiter whenever the meteorological conditions in Padua allowed him. As it became evident that the change in the configurations had visible changes on a scale of a few hours [^3] The period of Io is just about 1.8 days. Galileo decided to observed the system at different times during the same night. Such multiple observations occurred for example on Jan. 15 (id 7, 8), Jan. 17 (id 10, 11), Jan. 19 (id 13, 14), Jan. 20 (id 15, 16, 17), Jan. 22 (id 19, 20), Jan. 23 (id 21, 22), Jan. 24 (id 23, 24), Jan. 26 (id 26, 27), Jan. 31 (id 30, 31), Feb. 2 (id 33, 34), Feb. 4 (id 36, 37), Feb. 11 (id 43, 44, 45), Feb. 15 (id 48, 49, 50), Feb. 18 (id 53, 54), Feb. 26 (id 58, 59), Feb. 28 (id 62, 63). The range of these measurements taken during the same night spans up to about six hours. On Jan. 20 for example the first observation was done 1h15’ after sunset and the last 7h after sunset so at about 6 p.m. and 1 a.m.. Generally Galilei reported the configuration of the satellites only except in four occasions (observation 60, 61, 62, 64) when he also noted down the position of a nearby star [^4] We have identified this star as HD32811 or SAO 76962m or HIP 23784. It has mag. 7.16, a distance of 872.09 light years, type B8/9V and is located at 04h 43m 18.2s +21o549.9” according to the software Stellarium.. This was intended to clarify the motion of Jupiter and of its satellites with respect to the “really-fixed” stars. The satellites are marked as asterisks and the size of the symbol is a proxy for the relative brightness. In many occasions Galileo also documented the fact that the satellites appeared to be not completely aligned (see for example observations 6, 7, 11, 16, 17, 19, 23, 27, 30, 31). As we will see in the following when comparing with modern simulation also the recordings of these displacements from the ecliptic plane are remarkably accurate.
3 Digitization of the sketches

### v2

@@ Abstract
We present a new careful and comprehensive analysis the observations of the satellites of Jupiter from the Sidereus Nuncius that extends and complements previous similar studies. Each observation is compared to the predictions obtained using a modern sky simulator, verifying and trying to understand them individually. The work considers both the information that can be extracted from the sketches and the angular measurements reported by Galilei in the text. The consistency and relation between these two data-sets is explored. The angular measurements also allow assessing the absolute accuracy of the angular measurements in relation to modern ephemerides. Exploiting the simulations we use the data to evaluate the performances of the telescope in terms of separation power of close-by satellites and the inefficiency in the detection connected to the proximity to the disk. A sinusoidal fit of the data obtained from the available sketches, allows measuring the relative major semi-axes of the satellites’ orbits and their periods with a statistical precision of 2-4% and 0.1-0.3% respectively. The posterior fit error is used to estimate the accuracy in the reconstruction of the elongations. We show that with this data one can infer in a convincing way the third law of Kepler for the Jupiter system. The 1:2 and 1:4 orbital resonance between the periods of Io and Europa/Ganymede can be determined with % precision. In order to obtain these results it is important to separate the four datasets. This operation was an extremely difficult task for Galilei as the analysis will evidence. Nevertheless we show how some indication on the periods emerge from the using the modern Lomb-Scargle technique without having to separate the four data-sets. We briefly extend the use of the simulator to verify the accuracy in the seven observations of the Moon and the performance in reproducing the Pleiades, the Orion belt, the Orion head and the Beehive cluster. Finally we present some results on the achievable imaging of Jupiter and the Moon obtained with a replica of the instrument. This exercise highlights the challenges connected with these observations and confirms the excellence underlying this amazing set of early scientific data.
@@ 1 Introduction
[…]
@@
On January 8, Jupiter lies to the East of the three stars (second entry of Tab. 3). In reality the easternmost “star” (Callisto) was still there, to the East side, but Galilei does not notice it, this time. This might mean that indeed he was still not completely aware of what was going on and a still a bit less focused on these observations. As we will see he almost never missed a “detectable” satellite by mistake again thereafter. In particular Callisto will always be observed in the following two months, unless when it was too close to the disk. The three stars were then in reality not the same three satellites: Io had passed from the left to right side and Europa had become visible to the right after being too close to Jupiter the day before.
At this point, two intriguing observations was making Galilei eager to observe the situation in the following days: first the relative separations of the “stars” appear a little different; second, one had to assume that Jupiter was moving eastward (i.e. in retrograde motion) while in reality, as the opposition had already passed in 8 Dec. 1609, it was expected to be moving westward. The path of the Jupiter system in those days is shown in Fig. 1(data from [1]). The combined motion is quite complex. Galilei cross checks the motion of the planet with respect to a “truly” fixed star (HD32811) only much after with the last observations of February and early March. It is interesting to notice that the same star had appeared much closer to Jupiter as the planet was overcoming it from East to West. After the second day anyway the baseline, unexciting hypothesis that these were the same three fixed stars was still in place but a bit shaky already. This amazing story has been covered in the literature by historians of Physics and Astronomy in great detail. Of particular interest is the paper of 2011 by Gingerich and van Helden[2] where they elaborate in detail on how Galileo became convinced that the four points of light were satellites and not stars.
Figure 1: Motion of Jupiter (black) and of its satellites (Io in blue, Europa in red, Ganymede in green and Callisto in Orange) across the sky from Jan. 7 1610 to March 2 1610. The marker points the position of the star HD32811 that was recorded by Galileo in observations 61, 62 and 64 (see Tab. 3.)
@@
But coming back to those days, the cloudy sky of Padua on the 9th creates some “suspense” but the scene of the 10th is a real breakthrough as Jupiter is now West of the two “stars” instead of even farther to the East: this is inconsistent with Jupiter “overcoming” the two fixed stars by moving East. Moreover one of the three stars had disappeared indicating that something more interesting was going on. A thrilling adventure that will accompany him uninterruptedly for each clear night for the following two months was just about to start. It is remarkable how many clear days there were in Padua (/Venice) during those winter months: just ten days were cloudy out of about sixty. On average nowadays the fraction of cloudy days is 40-50% according to [3]. This detailed account of the dance of the satellites of Jupiter is an history of scientific rigor and excitement: it lead him to observe until late at night in several occasions.
Earlier in the Sidereus Galilei encourages astronomers in determining the periods of the satellites as this was not possible to him due to the restrictions in time. Still, despite this consideration it is amazing how accurately Galilei decided to document what he saw. It is also interesting to notice how early Galilei wanted to focus on the determination of the periods, an activity that will keep him busy for years. Despite being insufficient to estimate the periods, according to Galilei, the set of observations is extremely rich and well professionally documented. It is kind of amazing considering that this was just meant as a quick first release in support of the discovery.
@@
In fact these observations were so exceptional and groundbreaking to deserve such a solid and complete basis of evidence despite this might seem redundant to our eyes. In [4] we find a good account of the trepidation with which Galilei was awaiting the reactions of his contemporaries colleagues and especially the one of Kepler.
@@
It should also be not forgotten that Galilei might have felt the pressure of releasing the book as soon as possible due to the risk of being anticipated by other scientists. This risk was indeed quite well founded given the large spread of low magnification/low quality telescopes that had interested Europe already since 1608[5, 4] and the actual presence of competitors (Mayr, Harriot [6]). Despite this he deemed necessary to report such detailed data until the very close to the publication deadlines. The dedication to Cosimo de Medici is dated 12 Mar. 1610 and on Mar. 13 the book was in press of Tommaso Baglioni in Venice. The last observation dates back to just ten days earlier. This choice tells a lot about of the mindset of Galilei and this emphasis on data cannot but induce a certain degree of admiration in modern experimentalists.
In this work we present a new, detailed analysis of the observations of the satellites of Jupiter reported by Galileo Galilei in the Sidereus Nuncius [7] during the period from January 7 to March 3, 1610 besides some checks of other observations found in the same book.
1.1 Previous results
The idea of comparing this exciting data-set to modern ephemeris is not new. This approach was pioneered in 1962-1964, by the Belgian astronomer Jean Meeus that re-examined Galileo’s observations of Jupiter’s moons [8, 9], showing that all four were visible on 7 Jan. 1610. He first showed how it was not possible to resolve satellites being too close to each other or closer than about three Jovian radii from the planet’s limb. Using his newly computed satellite tables, Meeus reconstructed their configurations from 1600 to 2200.
Later, the historian Stillman Drake extended Meeus’ work by analysing all of Galileo’s Jupiter observations[10, 11].
In 1982 John Roche[6] analyzed the observed elongations and compared them against modern ephemerides. He discusses how the measurements of Galilei were later used by Harriot and combined with his measurements to calculate the periods of the satellites. In doing so this he performs a careful evaluation of a subset of the data of the Sidereus showing something already noticed by North [12] i.e. that beside a random error there is a systematic component that makes the angular measurements overestimate the real elongations by large factors. Roche interestingly mentions (pag. 25) that this is also observed in the data of Mayr and Harriot. He reports an overestimation of 1.4 for Galilei, 1.3 for Harriot and 1.35 for Mayr. This point is also discussed in the book by Stillman Drake of 1983[13]. We will come back to this interesting point and the hypotheses of Roche later during the discussion of our results that fully support this observation with a much larger data-sample. We shows that it is a real scale factor that is consistent for the four satellites .
In his 1983 commentary on Sidereus Nuncius, Drake compared Galilei’s notes with modern ephemerides and concluded that Galilei could generally resolve two satellites only if they were separated by more than $\sim 10^{{}^{\prime\prime}}$ [13].
In 1997 Levi, F.A., Levi-Donati also presented a verification of the observations of the Sidereus [14] using several programs (Planet 4.0, Voyager 1.0, Microprojects Astro 1.3 and 1.4, StarryNight Mac 1.0.1). In 2012 Bernieri published a work[15] where he revisits Galileo’s observations of Jupiter’s moons, using a modern planetarium software (TheSky6) to reconstruct the actual angular separations of Jupiter’s satellites as Galileo would have seen them. He carefully accounts for uncertainties in Galileo’s timing ($\pm$15 min) and the software’s positional precision ($\sim 1^{\prime\prime}$). This analysis reveals that the telescope could only resolve satellite–satellite separations greater than approximately 19–20′′ (much larger than earlier estimates of $\sim$10′′ by Drake) and that satellite–planet separations needed to exceed roughly 50′′ for detection, due to glare and optical aberrations. Additionally, Bernieri calculates that the mean error in Galileo’s recorded angular measurements is about 57′′, matching Galileo’s own estimation in Sidereus. Notably, these measurement errors show a systematic positive bias and correlate with larger separations—indicating an effect that we will discuss in more detail in our analysis.
A study involving a sinusoidal fit of the orbits by Alessandro Bettini is presented in [16] (2016). In Fig. 4.20. Bettini considers later data taken in 1611 and performs a sinusoidal fit of the elongations of Callisto and Ganymede. He observes that Galilei obtained a very good agreement in the periods (1.76, 3.53, 7.16, 16.3 days to be compared with modern values of 1.77 3.55, 7.1, 16.75). Concerning the elongations the knowledge evolved as shown in Tab. 1.
Table 1: Angular radii of the orbits of Jupiter’s satellites (in units of Jupiter radii). From [16].
Epoch Io Europa Ganymede Callisto
1610 3.5 5.7 8.8 15.3
1611 3.8 6.2 8.4 15
1611? 4.0 7.0 10.0 15
1612 5.7 8.6 14.0 “almost 25”
Modern 5.58 8.88 14.16 24.90
Bettini also shows how these data allow to verify the third Kepler on the Jovian system and test the universality of the gravitational constant for the Jovian and the solar system.
Among the earlier work on the subject it is worth mentioning a very nice web page developed at the same time of Bernieri’s work by Ernie Wright [17] which provides animations of the satellites and a direct comparison for several configurations.
In this work we extend significantly the analysis of the Sidereus Nuncius dataset:
  • We present sinusoidal fits of the elongations vs time. From the fit parameters we show a comparison of orbital parameters with modern expectations. Data have been corrected from Italian to modern hours (Sec. 2.2). The residuals of the fit are used to estimate the resolution of the measurements.
  • We present a detailed comparison of two datasets (Sec. 22.1): the one coming from an analysis of the drawings in the published Sidereus and the angular data given by Galilei in the text, starting from Jan. 12. Using this dataset, we can, in principle, characterize in absolute terms the accuracy of the angular measurements by comparing with the ephemeris providing the expected elongations in degrees, and not anymore in terms of the Jupiter disk whose definition is affected by the systematics on its determination with the telescope. To check this hypothesis, we test if the angular data support the evidence of the decrease in apparent angle of the Jupiter system as it moved away from the Earth from January to March (a 15% effect). In addition this dataset is closer to the real measurements of Galilei as it does not pass through the additional layer of complexity introduced by the printing process.
  • For the first time we show how a frequency analysis of sparse data (Lomb-Scargle periodograms) allows getting a fairly solid evaluation of the orbital periods of Callisto and Ganymede without having to disentangle the individual satellites;
  • We extend the study proposed by Bettini to this dataset (test of the third Kepler law);
  • We introduce a check of the 1:2:4 resonance of the periods of the innermost satellites;
  • We evaluate the correctness of all the configurations and discuss the only known anomaly of February 12. Each configuration drawn by Galilei is set aside to the prediction from the software allowing to evaluate in a direct and intuitive way the reasons for missing observations;
  • We reevaluate the resolving power of pairs of satellites and the inefficiency introduced by the proximity to Jupiter’s disk by analyzing the distribution of the positions of the missed satellites or missed pairs at “truth-level” using the simulator;
  • We extend the comparison with the simulator for other observations reported in the Sidereus (the Pleiades, the Beehive cluster, the Orion belt and head) in Sec. 6. Here we also present some comparisons between the original notes and the Sidereus and a cross check of the absolute angular measurements of Galilei for a known stellar pattern;
  • We then consider the seven inkwashes of the Moon that were the basis for the engravings of the Sidereus and, using the generally accepted dates of these observations, that are widely reported in the previous literature, we accurately compare them to modern expectations of the Moon appearance with the support of the simulator.
  • In the last part of the paper we describe the technical difficulties involved in collecting these data by the direct experience obtained by observing Jupiter and the Moon with a replica of one of the telescopes of Galilei that we developed for this purpose; We give a particular emphasis on the difficulty of the observations introduced by the narrowness of the field of view, a critical aspect for the success of these observations, that nevertheless is generally not particularly considered.
2 Analysed data
This analysis focuses on the first observations presented in the Sidereus Nuncius without considering those performed in later years by Galileo himself. We do not consider here the information provided by the handwritten notes of Galilei that are carefully analysed in the literature: [18, 19].
@@ 2 Analysed data
The images were extracted from the Sidereus Nuncius dititally available at [7]. We show the collection of the sketches in chronological order with a progressive numeric labeling in Figs. 2-3.
@@
In total Galileo presents 65 sketches in 1610 from January 7th up to March 1st basically inspecting Jupiter whenever the meteorological conditions in Padua or Venice allowed him. Here we consider only 64 sketches because one is a duplicate of a Jupiter configuration with the addition of the above-mentioned fixed reference star. As it became evident that the change in the configurations had visible changes on a scale of a few hours [^1] The period of Io is just about 1.8 days. Galileo decided to observed the system at different times during the same night. Such multiple observations occurred for example on Jan. 15 (id 7, 8), Jan. 17 (id 10, 11), Jan. 19 (id 13, 14), Jan. 20 (id 15, 16, 17), Jan. 22 (id 19, 20), Jan. 23 (id 21, 22), Jan. 24 (id 23, 24), Jan. 26 (id 26, 27), Jan. 31 (id 30, 31), Feb. 2 (id 33, 34), Feb. 4 (id 36, 37), Feb. 11 (id 43, 44, 45), Feb. 15 (id 48, 49, 50), Feb. 18 (id 53, 54), Feb. 26 (id 58, 59), Feb. 28 (id 62, 63). The range of these measurements taken during the same night spans up to about six hours. On Jan. 20 for example the first observation was done 1h15’ after sunset and the last 7h after sunset so at about 6 p.m. and 1 a.m.. Generally Galilei reported the configuration of the satellites only except in four occasions (observation 60, 61, 62, 64) when he also noted down the position of a nearby star. This star is named HD32811 or SAO 76962m or HIP 23784 in modern catalogues. It has a visual magnitude of 7.16, a distance of 872.09 light years, type B8/9V and is located at 4h 43m 18.2s +21549.9′′. These data were obtained with the online tool Stellarium web that we will use extensively in the following[20, 21, 22]. The observation of HD32811 was intended to clarify the motion of Jupiter and of its satellites with respect to the “really-fixed” stars and came at end of the campaign. The satellites are marked as asterisks and the size of the symbol is a proxy for the relative brightness. Galilei always takes care of specifying if the satellites appeared as aligned or not. See for example observations 6, 7, 11, 16, 17, 19, 23, 27, 30, 31 in Appendix A. As we will see in the following the recordings of these displacements from the ecliptic plane are remarkably accurate and recorded with full awareness.
The observations are generally thought to have taken place in the garden of Galilei’s house in Padua[23], located just north of the Basilica del Santo and the Odeo Cornaro at 45.4032, 11.8818. From Jan. 30 to Feb. 12 we know from the correspondence that Galilei was in Venice instead[15]. The distribution of the position of Jupiter in alt-azimuthal coordinates is shown in Fig. 4 with color codes indicating the observation time during the night. It is apparent how Galilei had a rather clear horizon as the observations are almost uniformly distributed. The asymmetry towards the East is due to the fact that the observations could not have been made during daytime. In general Galilei used to observe immediately after the sunset. The late observations at about midnight are limited by the fact that Jupiter started to be low on the horizon. The altitude angle for the telescope ranged from about 30 to 70. It is interesting to notice that the highest point of the Basilica del Santo is 72 m at a distance of about 200 m has hence an altitude of about 20, it covers an azimuth between roughly 183 and 212. The Odeo and Loggia Cornaro were at about 50 m with an height of about 12 m thus subtending an angle of 13 in an azimuth window between 174 and 143. The house itself is oriented almost east-west so the late observations might have been limited by the house itself as they happened about 5 north of the West direction. This setting is hence perfectly consistent with the observations [^2] An early hypothesis that the tower of Porta Molino, along the medieval walls of Padua, was also the setting of the observation has been recently dismissed.
Figure 4: Distribution of the observations in azimuth and altitude for the supposed location of the observations in Padua (45.4032, 11.8818, currently at via Galileo Galilei 18). The color code indicates CET time of the observation.
2.1 Dataset from the digitization of the sketches